OpenStudy (gabbyalicorn):

What is the circumference of a circle with a radius of 1.4 m?

OpenStudy (gabbyalicorn):

I'm guessing 2 x 3.14 x 1.4 ?

OpenStudy (michele_laino):

yes! what do you get?

OpenStudy (gabbyalicorn):

8.792

OpenStudy (michele_laino):

here we have to apply this formula: \[C = 2 \times \pi \times 1.4 = ...inches\]

OpenStudy (gabbyalicorn):

I still have 8.792

OpenStudy (michele_laino):

no, please we have to do this computation: \[C = 2 \times \pi \times 1.4 = 2 \times 1.4\pi = 2.8 \times \pi \;inches\]

OpenStudy (michele_laino):

oops.. \[C = 2 \times \pi \times 1.4 = 2 \times 1.4\pi = 2.8 \times \pi \;meters\]

OpenStudy (gabbyalicorn):

C= 8.792=8.792 = 8.792

OpenStudy (michele_laino):

here you have to multiply 2*1.4 only, so our answer is: \[C = 2 \times \pi \times 1.4 = 2 \times 1.4\pi = 2.8 \times \pi \;meters\] and we have finished

OpenStudy (michele_laino):

\[C = 2.8 \times \pi \;meters\]

OpenStudy (gabbyalicorn):

2.8 x pi meters = 8.792 :I

OpenStudy (michele_laino):

you don't have to multiply 2.8*3.14

OpenStudy (gabbyalicorn):

okay, without pi it would be 2.8 which would be D ;D

OpenStudy (gabbyalicorn):

Thank you!

OpenStudy (michele_laino):

that's right! :)